OpenStudy (anonymous):

I need help solving the following basic linear equation: 1-3y/5=2

OpenStudy (anonymous):

is 3y or (1- 3y) over 5?

OpenStudy (anonymous):

\[1-3y \div5=2\]

OpenStudy (anonymous):

1-3y over 5 equals 2

OpenStudy (anonymous):

1-2=3y/5, -1*5=3y -5/3=y

OpenStudy (anonymous):

It says the answers is -3 that is why I am confused

OpenStudy (anonymous):

I got the same thing as you the first time

OpenStudy (anonymous):

(1-3y)/5=2, 1-3y=2*5 -3y=10-1, y=9/-3, y=-3

OpenStudy (anonymous):

@haganmc thank you it makes sense